# Maxitive functions with respect to general orders

### v2

@@ 4.1. Usual stochastic order
Let $\Phi_{\rm st}:=\{\phi_{u}\colon u\in(0,1)\}$ denote the family of functions $\phi_{u}\colon\mathcal{M}\to\mathbb{R}$ given by $\phi_{u}(\xi):=q_{\xi}(u)$. We identify $\Phi_{\rm st}$ with $(0,1)$, and endow $(0,1)$ with the Euclidean topology.444$(0,1)$ is not complete with the Euclidean distance. However, it is completely metrizable, that is, there exists an equivalent metric on $(0,1)$ such that $(0,1)$ with the relative Euclidean topology is a complete metric space space. In fact, every countable intersection of open subsets of a complete metric space is completely metrizable, see [27, Theorem 24.12]. Then it holds $\xi\preccurlyeq_{\rm st}\eta$ if and only if $\phi_{u}(\xi)\leq\phi_{u}(\eta)$ for all $u\in(0,1)$. Moreover, for every $\xi\in\mathcal{M}$, the mapping $u\mapsto\phi_{u}(\xi)$ is lower semicontinous. Therefore, the partially ordered space $(\mathcal{M},\preccurlyeq_{\rm st})$ satisfies the assumptions of Section 3.
@@
Let $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$ with $q^{\ast}(u):=\sup_{n\in\mathbb{N}}q_{\xi_{n}}(u)\leq\gamma(u)$ for all $u\in(0,1)$. Then the mapping $u\mapsto q^{\ast}(u)$ is real-valued, non-decreasing and left-continuous. Hence, there exists $\xi^{\ast}\in\mathcal{M}$ with $q^{\ast}(u)=q_{\xi^{\ast}}(u)$ for all $u\in(0,1)$. In particular, $\xi^{\ast}=\vee_{n\in\mathbb{N}}\xi_{n}$ and $\phi_{u}(\xi^{\ast})=q^{\ast}(u)\leq\gamma(u)$ for all $u\in(0,1)$. ∎
For a function $\psi\colon\mathcal{M}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}q_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$ for all $s\in\mathbb{R}$ and $u\in(0,1)$. If $\psi$ is countably maxitive, then the sublevel sets are countably directed, so that the mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. As a direct application of Theorem 3.1 and Lemma 2.3, we obtain the following representation result for maxitive functions w.r.t. the usual stochastic order.
Theorem 4.2.
Let $\psi\colon\mathcal{M}\to(-\infty,\infty]$ be a function. Then the following conditions are equivalent:
@@
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm st}$.
@@
    $\psi$ is monotone and has countably directed sublevel sets w.r.t. $\preccurlyeq_{\rm st}$.
@@
Let $\psi\colon\mathcal{M}\to(-\infty,\infty]$ be a function with the translation property. Then the following conditions are equivalent:
@@
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm st}$.
@@
Recall that the value at risk ${\rm VaR}_{u}\colon\mathcal{M}\to\mathbb{R}$ at level $u\in(0,1)$, which is defined by ${\rm VaR}_{u}(\xi):=q_{\xi}(u)$, is a monetary risk measure. A detailed discussion on risk measures can be found in [9]. The previous Corollary 4.3 is closely related to [26, Theorem 4] which provides a representation result for lower semicontinuous complete maxitive functionals $\mathcal{M}_{1}\to\mathbb{R}$ having the translation property. A related representation result was also shown in [1, Proposition 4] for the class of benchmark-loss VaR which were introduced by Bignozzi et al. [1].
@@
For every sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$ such that $\vee_{n\in\mathbb{N}}\xi_{n}$ exists, it holds ${\rm VaR}_{u}(\vee_{n\in\mathbb{N}}\xi_{n})=\sup_{n\in\mathbb{N}}{\rm VaR}_{u}(\xi_{n})$. Nevertheless, the function ${\rm VaR}_{u}\colon\mathcal{M}\to\mathbb{R}$ is not countably maxitive because it may happen that $\sup_{n\in\mathbb{N}}{\rm VaR}_{u}(\xi_{n})<\infty$ for an unbounded sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$. However, the function ${\rm VaR}_{u}$ can also be defined on the set $\mathcal{M}^{\ast}$ of all extended random variables $\xi\colon\Omega\to(-\infty,\infty]$. Then the supremum $\vee_{i\in I}\xi_{i}$ exists for every family $(\xi_{i})_{i\in I}\subset\mathcal{M}^{\ast}$ and ${\rm VaR}_{u}\colon\mathcal{M}^{\ast}\to(-\infty,\infty]$ is completely maxitive.
@@ 4.2. Increasing convex stochastic order
Let $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}_{1}$ with $\vee_{n\in\mathbb{N}}Q_{\xi_{n}}(u)\leq\gamma(u)$ for all $u\in(0,1)$. Since $\gamma$ is concave, it holds $Q^{\ast}(u):={\rm conc}(Q_{\xi_{1}},Q_{\xi_{2}},\cdots)(u)\leq\gamma(u)$ for all $u\in(0,1)$. Since $u\mapsto Q^{\ast}(u)$ is real-valued, we obtain $Q_{\xi^{\ast}}=Q^{\ast}$ where $\xi^{\ast}=\vee_{n\in\mathbb{N}}\xi_{n}\in\mathcal{M}_{1}$, and $Q_{\xi^{\ast}}(u)=Q^{\ast}(u)\leq\gamma(u)$ for all $u\in(0,1)$. ∎
For a function $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}_{1}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}Q_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$ for all $s\in\mathbb{R}$ and $u\in(0,1)$. If $\psi$ is countably maxitive, then the sublevel sets are countably directed, from which it is straightforward to verify that the mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued and concave for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. [^6] As for the concavity, we notice that for $\xi_{1},\xi_{2}\in A_{s}$, there exists $\xi^{\ast}\in A_{s}$ with $Q_{\xi^{\ast}}={\rm conc}(Q_{\xi_{1}},Q_{\xi_{2}})$ such that $\lambda Q_{\xi_{1}}(u_{1})+(1-\lambda)Q_{\xi_{2}}(u_{2})\leq Q_{\xi^{\ast}}(\lambda u_{1}+(1-\lambda)u_{2})$ for all $u_{1},u_{2},\lambda\in(0,1)$. As a direct application of Theorem 3.1 and Lemma 2.3, we obtain the following representation result for maxitive functions w.r.t. the increasing convex stochastic order.
Theorem 4.6.
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function. Then the following conditions are equivalent:
@@
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm icx}$.
@@
    $\psi$ is monotone and has countably directed sublevel sets w.r.t. $\preccurlyeq_{\rm icx}$.
@@
The expected shortfall is a convex risk measure, see e.g. [9] for further details. In particular, the functions ${\rm ES}_{u}$ have the translation property and also induce the partial order $\preccurlyeq_{\rm icx}$. However, by changing to the family $\Phi^{\prime}_{\rm icx}$, we also change the set of countably stable functions. More precisely, a function $\gamma\colon(0,1)\to[-\infty,\infty]$ is countably stable w.r.t. $\Phi_{\rm icx}$ if and only if the function $u\mapsto\frac{1}{1-u}\gamma(u)$ is countably stable w.r.t. $\Phi_{\rm icx}^{\prime}$. Moreover, the minimal penalty function $\alpha_{\min}^{\prime}$ w.r.t. $\Phi_{\rm icx}^{\prime}$ is given by
@@
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function with the translation property. Then the following conditions are equivalent:
@@
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm icx}$.
@@
The previous Corollary 4.7 is closely related to [26, Theorem 5]. Moreover, it is shown in [26] that the function ${\rm ES}_{u}$ does not always preserve the maximum of two randoms variables w.r.t. $\preccurlyeq_{\rm icx}$, i.e., the function ${\rm ES}_{u}$ is not maxitive w.r.t. $\preccurlyeq_{\rm icx}$.
@@ 4.3. Convex stochastic order
In particular, $(\mathcal{M}^{r}_{1},\preccurlyeq_{\rm cx})$ is a complete lattice. For $\psi\colon\mathcal{M}^{r}_{1}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}^{r}_{1}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}\colon\mathbb{R}\times(0,1)\to[-\infty,\infty]$ by $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}Q_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$. As a direct application of Theorem 3.1 and Lemma 2.3, we obtain the following representation result for maxitive functions w.r.t. the convex stochastic order.
Theorem 4.8.
Let $\psi\colon\mathcal{M}_{1}^{r}\to\mathbb{R}$ be a function. Then the following conditions are equivalent:
  • (i)
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm cx}$.
  • (ii)
    $\psi$ is monotone and has countably directed sublevel sets w.r.t. $\preccurlyeq_{\rm cx}$.
  • (iii)
    $\psi(\xi)=\min\left\{s\in\mathbb{R}\colon\sup_{u\in(0,1)}\{Q_{\xi}(u)-\alpha(s,u)\}\leq 0\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}^{r}.$
  • (iv)
    $\psi(\xi)=\sup_{u\in(0,1)}G(Q_{\xi}(u),u)\quad\mbox{ for all }\xi\in\mathcal{M}_{1}^{r},$
@@ 4.4. Increasing concave stochastic order
For a function $\psi\colon\mathcal{M}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}_{1}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}\bar{Q}_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$ for all $s\in\mathbb{R}$ and $u\in(0,1)$. If $\psi$ is countably maxitive, then the sublevel sets are countably directed, so that the mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. As a direct application of Theorem 3.1 and Lemma 2.3, we obtain the following representation result for maxitive functions w.r.t. the increasing concave stochastic order.
Theorem 4.10.
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function. Then the following conditions are equivalent.
@@
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm icv}$.
@@
    $\psi$ is monotone and has countably directed sublevel sets w.r.t. $\preccurlyeq_{\rm icv}$.
@@
where $\alpha_{\min}^{\prime}$ is the minimal penalty function w.r.t. $\Phi_{\rm icv}^{\prime}$. As a consequence of Corollary 3.3, we obtain the following result.
Corollary 4.11.
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function with the translation property. Then the following conditions are equivalent:
  • (i)
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm icv}$.
  • (ii)
    There exists a function $\alpha\colon(0,1)\to[-\infty,\infty)$ such that
    $\psi(\xi)=\sup_{u\in(0,1)}\left\{\overline{{\rm ES}}_{u}(\xi)-\tfrac{1}{1-u}\alpha(u)\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}.$
  • (iii)
    The mapping $u\mapsto\alpha_{\min}(0,u)$ is real-valued. Moreover, it holds
    $\psi(\xi)=\sup_{u\in(0,1)}\left\{\overline{{\rm ES}}_{u}(\xi)-\tfrac{1}{1-u}\alpha_{\min}(0,u)\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}.$

### v3

@@ 5.1. Usual stochastic order
Let $\Phi_{\rm st}:=\{\phi_{u}\colon u\in(0,1)\}$ denote the family of functions $\phi_{u}\colon\mathcal{M}\to\mathbb{R}$ given by $\phi_{u}(\xi):=q_{\xi}(u)$. We identify $\Phi_{\rm st}$ with $(0,1)$, and endow $(0,1)$ with the Euclidean topology.212121$(0,1)$ is not complete with the Euclidean distance. However, it is completely metrizable, that is, there exists an equivalent metric on $(0,1)$ such that $(0,1)$ with the relative Euclidean topology is a complete metric space space. In fact, every countable intersection of open subsets of a complete metric space is completely metrizable, see [33, Theorem 24.12]. Then, it holds $\xi\preccurlyeq_{\rm st}\eta$ if and only if $\phi_{u}(\xi)\leq\phi_{u}(\eta)$ for all $u\in(0,1)$. Moreover, for every $\xi\in\mathcal{M}$, the mapping $u\mapsto\phi_{u}(\xi)$ is lower semicontinuous. Therefore, the partially ordered space $(\mathcal{M},\preccurlyeq_{\rm st})$ satisfies the assumptions of Section 4.
@@
Let $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$ with $q^{\ast}(u):=\sup_{n\in\mathbb{N}}q_{\xi_{n}}(u)\leq\gamma(u)$ for all $u\in(0,1)$. Then, the mapping $u\mapsto q^{\ast}(u)$ is real-valued, non-decreasing and left-continuous. Hence, there exists $\xi^{\ast}\in\mathcal{M}$ with $q^{\ast}(u)=q_{\xi^{\ast}}(u)$ for all $u\in(0,1)$. In particular, $\xi^{\ast}=\vee_{n\in\mathbb{N}}\xi_{n}$ and $\phi_{u}(\xi^{\ast})=q^{\ast}(u)\leq\gamma(u)$ for all $u\in(0,1)$. ∎
For a function $\psi\colon\mathcal{M}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}q_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$ for all $s\in\mathbb{R}$ and $u\in(0,1)$. If $\psi$ is countably maxitive, then the sublevel sets are countably upwards directed, so that $u\mapsto\alpha_{\min}(s,u)$ is real-valued for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. As a direct application of Theorem 4.1 and Lemma 3.4, we obtain the following representation result for maxitive functions w.r.t. the usual stochastic order. A related representation result for maxitive functionals with respect to the usual stochastic order was investigated in parallel to our work in [4, Theorem 1].
Corollary 5.2.
Let $\psi\colon\mathcal{M}\to(-\infty,\infty]$ be a function. Then, the following conditions are equivalent:
@@
    $\psi$ is bounded countably maxitive w.r.t. $\preccurlyeq_{\rm st}$.
@@
    $\psi$ is monotone and has countably upwards directed sublevel sets w.r.t. $\preccurlyeq_{\rm st}$.
@@
Let $\psi\colon\mathcal{M}\to(-\infty,\infty]$ be a function with the translation property. Then, the following conditions are equivalent:
@@
    $\psi$ is bounded countably maxitive w.r.t. $\preccurlyeq_{\rm st}$.
@@
Recall that the value at risk ${\rm VaR}_{u}\colon\mathcal{M}\to\mathbb{R}$ at level $u\in(0,1)$, which is defined by ${\rm VaR}_{u}(\xi):=q_{\xi}(u)$, is a monetary risk measure; see [12]. The previous Corollary 5.3 is closely related to [32, Theorem 4] which provides a representation result for lower semicontinuous completely maxitive functions $\mathcal{M}_{1}\to\mathbb{R}$ having the translation property. A related representation result was also shown in [2, Proposition 4] for the class of benchmark-loss VaR, which were introduced by Bignozzi et al. [2].
@@
For every sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$ such that $\vee_{n\in\mathbb{N}}\xi_{n}$ exists, it holds ${\rm VaR}_{u}(\vee_{n\in\mathbb{N}}\xi_{n})=\sup_{n\in\mathbb{N}}{\rm VaR}_{u}(\xi_{n})$. Nevertheless, the function ${\rm VaR}_{u}\colon\mathcal{M}\to\mathbb{R}$ is not countably bounded maxitive because it may happen that $\sup_{n\in\mathbb{N}}{\rm VaR}_{u}(\xi_{n})<\infty$ for an unbounded sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$. However, ${\rm VaR}_{u}$ can also be defined on the set $\mathcal{M}^{\ast}$ of all extended random variables $\xi\colon\Omega\to(-\infty,\infty]$. Then, the supremum $\vee_{i\in I}\xi_{i}$ exists for every family $(\xi_{i})_{i\in I}\subset\mathcal{M}^{\ast}$ and ${\rm VaR}_{u}\colon\mathcal{M}^{\ast}\to(-\infty,\infty]$ is completely maxitive.
@@ 5.2. Increasing convex stochastic order
Let $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}_{1}$ with $\vee_{n\in\mathbb{N}}Q_{\xi_{n}}(u)\leq\gamma(u)$ for all $u\in(0,1)$. Since $\gamma$ is concave, it holds $Q^{\ast}(u):={\rm conc}(Q_{\xi_{1}},Q_{\xi_{2}},\cdots)(u)\leq\gamma(u)$ for all $u\in(0,1)$. Since $u\mapsto Q^{\ast}(u)$ is real-valued, we obtain $Q_{\xi^{\ast}}=Q^{\ast}$, where $\xi^{\ast}=\vee_{n\in\mathbb{N}}\xi_{n}\in\mathcal{M}_{1}$, and $Q_{\xi^{\ast}}(u)=Q^{\ast}(u)\leq\gamma(u)$ for all $u\in(0,1)$. ∎
For a function $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}_{1}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}Q_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$ for all $s\in\mathbb{R}$ and $u\in(0,1)$. If $\psi$ is countably maxitive, then the sublevel sets are countably upwards directed, from which it is straightforward to verify that the mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued and concave for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. [^23] As for the concavity, we notice that for $\xi_{1},\xi_{2}\in A_{s}$, there exists $\xi^{\ast}\in A_{s}$ with $Q_{\xi^{\ast}}={\rm conc}(Q_{\xi_{1}},Q_{\xi_{2}})$ such that $\lambda Q_{\xi_{1}}(u_{1})+(1-\lambda)Q_{\xi_{2}}(u_{2})\leq Q_{\xi^{\ast}}(\lambda u_{1}+(1-\lambda)u_{2})$ for all $u_{1},u_{2},\lambda\in(0,1)$. As a direct application of Theorem 4.1 and Lemma 3.4, we obtain the following representation result for maxitive functions w.r.t. the increasing convex stochastic order.
Corollary 5.6.
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function. Then, the following conditions are equivalent:
@@
    $\psi$ is bounded countably maxitive w.r.t. $\preccurlyeq_{\rm icx}$.
@@
    $\psi$ is monotone and has countably upwards directed sublevel sets w.r.t. $\preccurlyeq_{\rm icx}$.
@@
The expected shortfall is a convex risk measure; see [12]. In particular, the functions ${\rm ES}_{u}$ have the translation property and also induce the partial order $\preccurlyeq_{\rm icx}$. However, by changing to the family $\Phi^{\prime}_{\rm icx}$, we also change the set of countably stable functions. More precisely, a function $\gamma\colon(0,1)\to[-\infty,\infty]$ is countably stable w.r.t. $\Phi_{\rm icx}$ if and only if the function $u\mapsto\frac{1}{1-u}\gamma(u)$ is countably stable w.r.t. $\Phi_{\rm icx}^{\prime}$. Moreover, the minimal penalty function $\alpha_{\min}^{\prime}$ w.r.t. $\Phi_{\rm icx}^{\prime}$ is given by
@@
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function with the translation property. Then, the following conditions are equivalent:
@@
    $\psi$ is bounded countably maxitive w.r.t. $\preccurlyeq_{\rm icx}$.
@@
The previous Corollary 5.7 is closely related to [32, Theorem 5]. Moreover, it is shown in [32] that the function ${\rm ES}_{u}$ does not always preserve the maximum of two randoms variables w.r.t. $\preccurlyeq_{\rm icx}$, i.e., the function ${\rm ES}_{u}$ is not maxitive w.r.t. $\preccurlyeq_{\rm icx}$.
@@ 5.3. Convex stochastic order
In particular, $(\mathcal{M}^{r}_{1},\preccurlyeq_{\rm cx})$ is a complete lattice. For $\psi\colon\mathcal{M}^{r}_{1}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}^{r}_{1}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}\colon\mathbb{R}\times(0,1)\to[-\infty,\infty]$ by $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}Q_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$. As a direct application of Theorem 4.1 and Lemma 3.4, we conclude that a function $\psi\colon\mathcal{M}_{1}^{r}\to\mathbb{R}$ is countably maxitive with respect to $\preccurlyeq_{\rm cx}$, if it admits a representation similar to that in Corollary 5.6, with $\mathcal{M}_{1}$ replaced by $\mathcal{M}_{1}^{r}$.
@@ 5.4. Increasing concave stochastic order
For a function $\psi\colon\mathcal{M}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}_{1}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}\bar{Q}_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$ for all $s\in\mathbb{R}$ and $u\in(0,1)$. If $\psi$ is countably maxitive, then the sublevel sets are countably upwards directed, so that $u\mapsto\alpha_{\min}(s,u)$ is real-valued for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. As a direct application of Theorem 4.1 and Lemma 3.4, we obtain the following representation result for maxitive functions w.r.t. the increasing concave stochastic order.
Corollary 5.9.
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function. Then, the following conditions are equivalent.
@@
    $\psi$ is bounded countably maxitive w.r.t. $\preccurlyeq_{\rm icv}$.
@@
    $\psi$ is monotone and has countably upwards directed sublevel sets w.r.t. $\preccurlyeq_{\rm icv}$.
@@
where $\alpha_{\min}^{\prime}$ is the minimal penalty function w.r.t. $\Phi_{\rm icv}^{\prime}$. As a consequence of Corollary 4.4, we conclude that a function $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ with the translation property is countably maxitive with respect to $\preccurlyeq_{\rm icv}$, if it admits a representation similar to that in Corollary 5.7, with ${\rm ES}_{u}$ replaced by $\overline{{\rm ES}}_{u}$.